OpenStudy (anonymous):

5x + 2y = 3 (equation 1) 2x – 3y = 1 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

Show that the solution to the system of equations 7x – y = 4 and 2x – 3y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 2x + 5y = 3 and 3x – 2y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 9x + 4y = 5 and 7x – y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations –4x + 9y = 5 and 2x – 3y = 1 is the same as the solution to the given system of equations

OpenStudy (solomonzelman):

these are the 4 options, right?

OpenStudy (solomonzelman):

(if so, then ....) You know that you must keep the 2nd equation unchanged, and 1 of the options has the second equation typed incorrectly, or another option that doesn't have a second equation in it at all. These options that lack the correct 2nd equation can be eliminated right away.